OpenStudy (anonymous):

If a cylindrical tank is 50 ft high and the radius of the base is 9 ft, how much water will it hold?

OpenStudy (anonymous):

Volume = \[\pi \times r ^{2}\times h = \pi \times 9^{2}\times50 = 4050\pi\] cubic feet

OpenStudy (anonymous):

thank you!